OpenStudy (anonymous):

2 square root of 44 x cubed - square root of 7 - 99 x cubed + square root of 63

OpenStudy (anonymous):

What should you do?? Anyway represent it better tough to understand.

OpenStudy (anonymous):

\[\Large 2\sqrt {44{x^3}} - \sqrt 7 - 99{x^3} + \sqrt {63} \]

OpenStudy (anonymous):

Brought to you by MathType.

OpenStudy (anonymous):

Yeah but what is the question?

OpenStudy (anonymous):

I don't know, simplify maybe? \[\Large \begin{array}{l} 2\sqrt {44{x^3}} - \sqrt 7 - 99{x^3} + \sqrt {63} \\ = 2 \cdot 2\sqrt {11} \cdot \sqrt {{x^3}} - \sqrt 7 - 99{x^3} + 3\sqrt 7 \\ = 4\sqrt {11} \cdot \sqrt {{x^3}} + 2\sqrt 7 - 99{x^3} \end{array}\]

hero (hero):

By the way, what's cherry pot?

OpenStudy (anonymous):

Yeah Hero i was wondering that too :D